OpenStudy (anonymous):

Find the exact value of cos 60+ sin 30+ -tan 45

OpenStudy (anonymous):

Ask yourself these questions: cos 60º = ? sin30º = ? tan45º = ?

OpenStudy (anonymous):

.5+.5-1=0

OpenStudy (anonymous):

Exactly. Do you understand?

OpenStudy (anonymous):

yea but i thought exact value meant something in fraction form

OpenStudy (anonymous):

can you help me with something else?

OpenStudy (anonymous):

Well, in this case:\[\Huge \frac{1}{2}+\frac{1}{2} - 1 = 0\]Still the same.

OpenStudy (anonymous):

oh i see!

OpenStudy (anonymous):

In a right triangle, CT=6, CA=3 and m< CAT=90. what is the value of cotC?

OpenStudy (anonymous):